OpenStudy (anonymous):

Adena has a number cube whose sides are labeled 1 through 6. She rolls the number cube three times. What is the probability that Adena will roll 2, then 3, and then 4? IT HAS TO BE A FRACTION!!!! I WILL GIVE MEDAL AND FAN!

OpenStudy (anonymous):

@UnkleRhaukus @dan815 @Bahurverse_Luver @loopgirl @chmvijay @nincompoop @ErGTA123 @TIvan @Ziferot

OpenStudy (anonymous):

plz help, i am desperate!

OpenStudy (kropot72):

On the first roll, the probability of rolling a 2 is 1/6. On the second roll the probability of rolling a 3 is also 1/6, and the probability of rolling a 4 on the third roll is, you guessed it, 1/6. Therefore the required probability is given by: \[\large \frac{1}{6}\times\frac{1}{6}\times\frac{1}{6}=you\ can\ calculate\]

OpenStudy (anonymous):

okay, i will tel u what i get in a sec

OpenStudy (anonymous):

okay, i got 1/216

OpenStudy (kropot72):

You are correct.

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

wow, i should have known :P

OpenStudy (kropot72):

Sure is!

OpenStudy (anonymous):

thx

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

there is ur medal :)

OpenStudy (kropot72):

Thank you kindly.